# An Empirical Study of Kotlin-Java Interactions

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. According to recent studies, a significant of Android apps have been continuously migrated from Java to Kotlin (Ardito et al., 2020; Mateus and Martinez, 2020). The interoperability of Java and Kotlin’s design nature allows them to coexist and evolve smoothly in a single project. Meanwhile, understanding the dependencies among software entities is fundamental to various aspects of software analysis, such as architecture recovery, code smell detection, and software quality evaluation (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Extracting and understanding the dependency interactions between Java and Kotlin code is essential for conducting software analysis in such cross-language software systems. Despite the increasing popularity of Kotlin and its interoperability with Java, there is limited research on how Java and Kotlin interact with and depend on each other in real-world projects. Furthermore, recent studies (Li et al., 2023b) have shown that a large percentage of bug fixes in deep learning frameworks implemented in Python and C++ involve modifying source code in both programming languages. Moreover, the complexity of code changes for such bug fixes is significantly higher than that of fixes in single-programming-language scenarios (Li et al., 2023a). Similarly, many Android applications are developed using both Java and Kotlin, mirroring the dual-language nature of deep learning frameworks which commonly employ Python and C++. Unlike the interactions between Python and C++, which rely on their compiled public APIs, Java and Kotlin, both being based on the Java Virtual Machine (JVM), can extensively depend on each other at the source code level. This raises the question of whether Kotlin-Java software systems also incur higher maintenance costs and are prone to cross-language bugs, similar to other multi-language systems like Python-C++. In other words, what challenges do developers face in the implementation and maintenance of Kotlin-Java systems? Understanding these challenges can assist researchers and practitioners in designing solutions and tools to navigate these issues and ensure the development of high-quality Kotlin-Java software systems.
@@
  • We demonstrated that the source files involved in Kotlin-Java interactions require undergoing more commits and changes than those only involved in a single language. It proves that the Kotlin-Java interface is not stable and actively changing.
@@ 2. Approach
In this section, we describe our approach to extract Kotlin-Java dependencies and to identify recurring patterns in Kotlin-Java problematic interactions which incur maintenance cost.
@@ 2.1. Kotlin-Java Dependency Extraction
We implemented a tool called DependExtractor to extract dependencies between Kotlin and Java code entities. DependExtractor first utilized the Kotlin compiler to resolve Kotlin and Java source files into PSI (Program Structure Interface, similar to Abstract Syntax Tree) elements, including their types, reference types, and locations. Subsequently, it extracted dependency relations between pairs of PSI elements. Finally, these dependencies into a graph containing all PSI elements are aggregated and can be exported in JSON format, with each node representing a Kotlin or Java source file and each edge representing the dependencies between two source files. Following the definitions of dependency relations from related works (Jin et al., 2019; dep, [n. d.]; und, [n. d.]), DependExtractor is capable of extracting 9 types of Java-Java dependencies, 14 types of Kotlin-Kotlin dependencies, and 11 types of Kotlin-Java dependencies. As this paper focuses mainly on Kotlin-Java interactions, we list the 11 types of Kotlin-Java dependencies and their descriptions in Table 1.
Table 1. Kotlin-Java Dependency Types Extracted by DependExtractor
@@
Access an expression accesses a field or property (kotlin side)
@@
2.2. Common Mistake Identification in Kotlin-Java Interaction
To detect recurring patterns associated with Kotlin-Java problematic interactions, we initially mined code repositories to narrow the problematic commit scope for manual inspection. Subsequently, we conducted a thorough manual inspection of code changes to identify and classify code patterns related to Kotlin-Java problematic interactions.
@@ 3.1. Subjects
Since GitHub does not support searching repositories by multiple languages, we clicked and manually inspected each repository to assess the ratio of programming languages and to fulfill Criterion 1). We established a minimum threshold of 3% for Java code to ensure its active participation in the projects’ functionalities. Criteria 2) and 3) focus on selecting repositories that are actively maintained, which are essential for identifying relevant challenges. By applying these criteria, we identified 23 projects, and their demographic information is presented in Table 2.
@@ 3.2. Research Questions
RQ1. How do dependencies among code entities look like in Kotlin-Java projects? Kotlin is designed to be fully interoperable with Java, allowing code entities to be seamlessly accessed across both languages. Ideally, dependencies between Kotlin and Java should exhibit the same characteristics as those within homogeneous Kotlin or Java codebases. The answer to this question can enhance our understanding of such cross-language software systems and form the basis for further software analysis in Kotlin-Java projects.
RQ2. Is the maintenance cost in files involved in Kotlin-Java interactions significantly higher than in files without Kotlin-Java interactions? Recent studies on multi-language systems, such as those involving Python and C++, have shown that to address bugs occurring in the interface of two languages, developers need to modify code in both languages, resulting in a higher number of lines of code (LOCs) (Li et al., 2023a). This observation raises the question of whether Kotlin-Java software systems are also prone to such cross-language maintenance costs. The answer to this question could pinpoint maintenance hotspots within these Kotlin-Java systems.
RQ3. Are there common mistakes in Kotlin-Java interactions? Previous research proposed various techniques for detecting code smells or anti-patterns within a single-language codebase. More recent studies have found that bugs and security problems in multiple language systems (Li et al., 2023b). For instance, a recent study has uncovered security issues arising from interactions between Java and C++ code (Li et al., 2022). By investigating whether there exist common mistakes in Kotlin-Java interactions, we aim to evaluate the problems associated with Kotlin-Java interactions and pinpoint the scenarios that may require closer attention when handling such interactions.
@@ 4.1. Dependencies in Kotlin-Java Projects
Using the approach detailed in Section 2.1, we developed a tool capable of extracting dependencies within Kotlin-Java projects. We utilized this tool on the most recent versions of our 23 selected projects, and the results are presented in Table 3. This table includes columns labeled #J and #K, which denote the counts of Java and Kotlin source files, respectively. The J-J and K-K columns represent the number of dependencies between pairs of Java source files and pairs of Kotlin source files, respectively. Additionally, the K-J column indicates the number of dependencies from a Kotlin source file to a Java source file. It is interesting that in these 23 projects we did not observe any Java-Kotlin dependencies (DependExtractor was tested thoroughly). This may be due to our subject selection criteria, and the primary language of these projects is Kotlin, in which developers tend to use Kotlin to depend on Java for functionalities. All dependencies in this table are counted without weights. The last column shows the ratio of Kotlin-Java dependencies to the minimum of Java-Java and Kotlin-Kotlin dependencies. Since source files with fewer numbers can determine dependency intensity, we compare Kotlin-Java dependencies with the minimum dependencies in a single language to check the prevalence of their interactions.
Table 3. Dependency Statistics in Kotlin-Java Projects
@@
Table 4. Dependency Types in Kotlin-Java Interactions
@@
Figure LABEL:fig:deps depicts the circular layout of dependencies in three projects. The nodes on the circle’s edge represent source files. The blue, red, and purple curved lines connecting two nodes represent Kotlin-Kotlin, Java-Java, and Kotlin-Java dependencies, respectively. It is evident that Kotlin-Java dependencies are not confined to a specific area; instead, they are spread across the entire systems, proving Kotlin-Java interoperability in real-world software projects.
@@
To answer RQ1, the dependency extraction from 23 open-source Kotlin-Java projects spanning various domains reveals that Kotlin-Java dependencies and interactions are widespread. Kotlin-Java dependencies can occur more frequently than minor single-language interactions, with “Access” and “Call” being the top two most frequent dependency types. Furthermore, Kotlin-Java dependencies are distributed across the entire system rather than confined to a local area, thus validating the interoperability between Kotlin and Java.
@@ 4.2. Maintenance Cost in Kotlin-Java Projects
Recent studies on multi-language systems have revealed that the interface between two languages is vulnerable and susceptible to bugs, often resulting in increased maintenance costs for fixing these issues (Li et al., 2023a, 2022, b). This raises the question of whether the interface between Kotlin and Java faces similar challenges. We investigated this question to gain a better understanding of Kotlin-Java interactions and the associated challenges.
For each project, utilizing the dependency information got in RQ1 (see Section 4.1), we extracted the paths of Kotlin or Java source files involved in Kotlin-Java dependencies, denoted as Paths(kt/java). Subsequently, we traversed each project’s git commits and matched each modified file path of a commit with Paths(kt/java). If a match was found, we incremented #commit(aFile) by 1 and added the added and deleted lines of code (LOC) to #LOC(aFile). Similarly, we calculated #commit and #LOC for Java or Kotlin source files participating only in Java-Java or Kotlin-Kotlin interactions.
@@
Comparing the first column with the second-to-last column, we observe that in 17 out of 23 projects (highlighted in blue), the average number of commits for Kotlin source files participating in Kotlin-Java interactions is significantly larger than for Kotlin source files not participating in Kotlin-Java interactions. The number in the parentheses denotes the ratio of these two commit counts. For example, in the husi project, the average number of commits (23.7) for Kotlin source files participating in Kotlin-Java interactions is 6.1 times of that (3.9) for Kotlin source files not participating in Kotlin-Java interactions. On average across all 23 projects, Kotlin source files participating in Kotlin-Java interactions have 2.5 times more commits than Kotlin source files not participating in Kotlin-Java interactions.
However, as observed from Table 5, the LOC for Kotlin source files participating in Kotlin-Java interactions show no significant difference from Kotlin source files not participating in Kotlin-Java interactions. Only in these four projects (highlighted in blue in the last column) Kotlin source files participating in Kotlin-Java interactions have more LOCs on average. Furthermore, we conducted a t-test for these columns, and the t-test results between column #Avg_Cmt_K and #Avg_Cmt_Kj indicate a statistically significant difference at a 95% confidence level (df = 22, p-value = 0.016). Such results support the alternative hypothesis that #Avg_Cmt_Kj is significantly greater than #Avg_Cmt_K. However, the t-test results indicate no statistically significant difference between #Avg_LOC_K and #Avg_LOC_Kj (df = 22, p-value = 0.125). We conjecture this is because Kotlin source files participating in the interactions are often committed to modify only a small amount of code related to the facade, such as method calls, field access, parameter changes, etc.
To answer RQ2, the average number of commits and LOC for 23 open-source Kotlin-Java projects indicate that Kotlin source files participating in Kotlin-Java interactions tend to have more commits but not more LOCs. On average across all 23 projects, Kotlin source files involved in Kotlin-Java interactions have 2.5 times more commits compared to Kotlin source files that do not participate in Kotlin-Java interactions.
@@ 4.3. Common Mistakes in Kotlin-Java Interactions
Table 6 presents the data in the process of locating commits involving problematic Kotlin-Java interactions. Initially, there were 83,792 commits among these projects. After filtering commits with comments containing the words ’fix’, ’bug’, ’issue’ or ’exception’, there were 5,137 remaining commits. These 5,137 commits were compared with the dependency graphs of the latest version, resulting in 1,314 commits with modified files participating in Kotlin-Java interactions. Finally, we manually inspected these 1,314 commits and confirmed 103 commits containing fixes for problematic Kotlin-Java interactions. After discussion among three developers, we classified 98 of these 103 problematic Kotlin-Java interactions into 7 categories according to the issue and fixing strategies. Table 7 lists these categories and their occurrences. In the following, we provide an example for each of these 7 categories.

### v2

# Cross-Language Dependencies: An Empirical Study of Kotlin-Java

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM / IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM); October 20–25, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. According to recent studies, a significant of Android apps have been continuously migrated from Java to Kotlin (Ardito et al., 2020; Mateus and Martinez, 2020). The interoperability of Java and Kotlin’s design nature allows them to coexist and evolve smoothly in a single project. Meanwhile, understanding the dependencies among software entities is fundamental to various aspects of software analysis, such as architecture recovery, code smell detection, and software quality evaluation (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Extracting and understanding the dependency interactions between Java and Kotlin code is essential for conducting software analysis in such cross-language software systems. Despite the increasing popularity of Kotlin and its interoperability with Java, there is limited research on how Java and Kotlin interact with each other in real-world projects.
Furthermore, recent studies (Li et al., 2023b) have shown that a large percentage of bug fixes in deep learning frameworks implemented in Python and C++ involve modifying source code in both programming languages. Moreover, the complexity of code changes for such bug fixes is significantly higher than that of fixes in single-programming-language scenarios (Li et al., 2023a). Unlike the interactions between Python and C++, which rely on their compiled public APIs, Java and Kotlin, both being based on the Java Virtual Machine (JVM), can extensively depend on each other at the source code level. This raises the question of whether Kotlin-Java software systems also incur higher maintenance costs and are prone to cross-language bugs, similar to other multi-language systems like Python-C++. Moreover, objects passing through the Python-C interface can incur vulnerabilities due to the different rules in these two languages. For example, if handled incorrectly, a Python list passed to a C program can incur Buffer Overflow due to the fixed-size array in C (Li et al., 2022). We are curious about the challenges developers face in the implementation and maintenance of Kotlin-Java systems and whether similar bugs or vulnerabilities exist in the Kotlin-Java interface. Understanding these challenges can assist researchers and practitioners in designing solutions and tools to navigate these issues and ensure the development of high-quality Kotlin-Java systems.
@@
  • We demonstrated that the source files involved in Kotlin-Java interactions require undergoing more commits than those only involved in a single language. It also proves that the Kotlin-Java interface is actively changing.
@@
Data Availability Statement: The data and implemented tool can be accessed through this link (Anonymous, 2024).
@@ 2.1. Kotlin-Java Dependency Extraction
DependExtractor Tool: We implemented a tool called DependExtractor in pure Kotlin to extract dependencies between Kotlin and Java code entities. DependExtractor first utilized the Kotlin compiler to resolve Kotlin and Java source files into PSI (Program Structure Interface, similar to Abstract Syntax Tree) elements, including their types, reference types, and locations. Subsequently, it extracted dependency relations between pairs of PSI elements. Finally, these dependencies into a graph containing all PSI elements are aggregated and can be exported in JSON format, with each node representing a Kotlin or Java source file and each edge representing the dependencies between two source files. Following the definitions of dependency relations from related works (Jin et al., 2019; dep, [n. d.]; und, [n. d.]), DependExtractor is capable of extracting 9 types of Java-Java dependencies, 14 types of Kotlin-Kotlin dependencies, and 11 types of Kotlin-Java dependencies. As this paper focuses mainly on Kotlin-Java interactions, we list the 11 types of Kotlin-Java (K-J) or Java-Kotlin (J-K) dependencies and their descriptions in Table 1.
Table 1. Supported K-J or J-K Dependency Types
@@
Access an expression accesses a Kotlin/Java field or Kotlin property
@@
PT and Extension occur from Kotlin entities to Java entities; Other DepTypes can occur either from Kotlin to Java or from Java to Kotlin.
DependExtractor Accuracy Validation: Besides a large amount of unit testing, we also used a Kotlin-Java project with verified dependencies as the ground truth to conduct an integrated test. This ground truth project (implemented in 108 Java and 22 Kotlin source files) is our pre-release version of DependExtractor. Just like software systems which migrate from Java to Kotlin, our tool also went through this process. As we are already familiar with the Kotlin and Java source files in this project, we feel confident to inspect the dependencies of it. In addition, this project contains 5,424 dependency instances, including all 11 dependency types in K-J or J-K, which is a good candidate to test DependExtractor’s ability to extract various dependency instances. Specifically, two authors, each with 4 years of Java and 3 years of Kotlin development experience, traced entities line by line in each source file to list the dependency instances independently. After completing this, they discussed the results until an agreement was reached. This manual inspection served as ground truth and were compared with the JSON result generated by Depends-Kotlin. The comparison shows that our tool can correctly capture 5,387 out of 5,424 dependencies (98.02% precision and 99.34% recall), increasing our confidence in applying DependExtractor in future experiments. Due to space limitations, the inspected ground truth and the results from the DependExtractor tool are shared in the provided Zenodo link (Anonymous, 2024).
2.2. Common Mistake Identification in Kotlin-Java Interaction
To identify Kotlin-Java problematic interactions, we initially mined code repositories to narrow the problematic commit scope for manual inspection. Subsequently, we conducted a thorough manual inspection of code changes to identify and classify code patterns related to Kotlin-Java problematic interactions.
@@ 3.1. Subjects
Since GitHub does not support searching repositories by multiple languages, we clicked and manually inspected each repository to assess the ratio of programming languages and to fulfill Criterion 1). We established a minimum threshold of 3% for Java code to ensure its active participation in the projects’ functionalities, not just some random Java code. Therefore, we chose a 3% threshold for Java code, rather than 1% or 2%. Criteria 2) and 3) focus on selecting repositories that are actively maintained, which are essential for identifying relevant challenges. By applying these criteria, we identified 23 projects, and their demographic information is presented in Table 2.
@@ 3.2. Research Questions
RQ1. Are dependencies between Kotlin and Java prevalent in Kotlin-Java projects? Kotlin is designed to be fully interoperable with Java, allowing code entities to be seamlessly accessed across both languages. Ideally, dependencies between Kotlin and Java should exhibit the same characteristics as those within homogeneous Kotlin or Java codebases. The answer to this question can enhance our understanding of such cross-language software systems.
RQ2. Do files involved in Kotlin-Java interactions cost more commits and lines of code (LOCs) than files without Kotlin-Java interactions? Recent studies on multi-language systems, such as those involving Python and C++, have shown that, to address bugs occurring in the interface of two languages, developers need to modify code in the interface, resulting in a higher number of lines of code (LOCs) (Li et al., 2023a). This observation raises the question of whether Kotlin-Java software systems are also prone to such cross-language maintenance activities. The answer to this question could pinpoint maintenance hotspots within these Kotlin-Java systems.
RQ3. Are there common mistakes in Kotlin-Java interactions? Previous research proposed various techniques for detecting code smells or anti-patterns within a single-language codebase. More recent studies have found that bugs and security problems in multiple language systems (Li et al., 2023b). For instance, a recent study has uncovered security issues arising from interactions between Java and C code (Li et al., 2022). By investigating whether there exist common mistakes in Kotlin-Java interactions, we aim to evaluate the problems associated with Kotlin-Java interactions and pinpoint the scenarios that may require closer attention when handling such interactions.
@@ 4.1. Dependencies in Kotlin-Java Projects
We applied DependExtractor on the most recent versions of our 23 selected projects, and the results are presented in Table 3. It is interesting that in these 23 projects we did not observe any Java-Kotlin dependencies (DependExtractor was tested thoroughly). This may be due to our subject selection criteria, and the primary language of these projects is Kotlin, in which developers tend to use Kotlin to depend on Java for functionalities. All dependencies in this table are counted without weights. The last column shows the ratio of Kotlin-Java dependencies to the minimum of Java-Java and Kotlin-Kotlin dependencies. Similar to the Barrel principle, if a project contains only a few Java source files, the number of cross-language dependencies should not be disproportionately large. Therefore, the smaller number of source files determines the cross-language dependencies. We compared Kotlin-Java dependencies with the minimum of J-J and K-K to check whether cross-language dependencies are more prevalent than single-language dependencies.
Table 3. Overall Dependencies in the Subjects
@@
#J and #K denote the number of Java and Kotlin files; J-J, K-K and K-J denote the dependencies number from Java to Java, Kotlin to Kotlin, and Kotlin to Java entities, respectively.
@@
Figure 1. Circular Layout of Dependencies
Table 4. Statistics of Dependencies from Kotlin to Java entities (K-J) in the subjects
@@
Figure 1 depicts the circular layout of dependencies in three projects. The nodes on the circle’s edge represent source files. The blue, red, and purple curved lines connecting two nodes represent Kotlin-Kotlin, Java-Java, and Kotlin-Java dependencies, respectively. It is evident that in these three projects, Kotlin-Java dependencies are not confined to a specific area; instead, they are spread across the entire systems. We observed this in 17 out of 23 projects, except for the Flap, grpc-kotlin, intellij-kotlin, karibu-testing, legado, and LiquidBounce projects. One explanation is that the migration process, legacy code, or design in these systems differs from the other 17 systems. We will investigate these observations in the future.
@@
To answer RQ1, the dependency extraction from 23 open-source Kotlin-Java projects spanning various domains reveals that Kotlin-Java interactions are widespread, thus validating the interoperability between Kotlin and Java. Kotlin-Java dependencies can occur more frequently than minor single-language interactions, with “Access” and “Call” being the top two most frequent dependency types. Furthermore, in most projects Kotlin-Java dependencies are distributed across the entire system rather than confined to a local area.
@@
#Avg_Cmt_K and Avg_LOC_K denote the average commit number and LOC for Kotlin files only involving K-K dependencies;
#Avg_Cmt_J and Avg_LOC_J denote the average commit number and LOC for Kotlin files only involving J-J dependencies;
#Avg_Cmt_Kj and Avg_LOC_Kj denote the average commit number and LOC for Kotlin files participating K-J dependencies.
@@ 4.2. Maintenance Cost in Kotlin-Java Projects
Recent studies on multi-language systems have revealed that the interface between two languages is vulnerable and susceptible to bugs, often resulting in increased maintenance costs for fixing these issues (Li et al., 2023a, 2022, b). This raises the question of whether the interface between Kotlin and Java faces similar challenges.
This research question aims to gain a better understanding of Kotlin-Java interactions. For each project, utilizing the dependency information got in RQ1 (see Section 4.1), we extracted the paths of Kotlin or Java source files involved in Kotlin-Java dependencies, denoted as Paths(kt/java). Subsequently, we traversed each project’s git commits and matched each modified file path of a commit with Paths(kt/java). If a match was found, we incremented #commit(aFile) by 1 and added the modified lines of code (including insertions and deletions) to #LOC(aFile). Similarly, we calculated #commit and #LOC for Java or Kotlin source files participating only in Java-Java or Kotlin-Kotlin interactions.
@@
Comparing the first column with the second-to-last column, we observe that in 17 out of 23 projects (highlighted in orange), the average number of commits for Kotlin source files participating in Kotlin-Java interactions is significantly larger than for Kotlin source files not participating in Kotlin-Java interactions. The number in the parentheses denotes the ratio of these two commit counts. For example, in the husi project, the average number of commits (23.7) for Kotlin source files participating in Kotlin-Java interactions is 6.1 times of that (3.9) for Kotlin source files not participating in Kotlin-Java interactions. On average across all 23 projects, Kotlin source files participating in Kotlin-Java interactions have 2.5 times more commits than Kotlin source files not participating in Kotlin-Java interactions.
However, as observed from Table 5, the LOC for Kotlin source files participating in Kotlin-Java interactions show no significant difference from Kotlin source files not participating in Kotlin-Java interactions. Only in these four projects (highlighted in orange in the last column) Kotlin source files participating in Kotlin-Java interactions have more LOCs on average. Furthermore, we conducted a t-test for these columns, and the t-test results between column #Avg_Cmt_K and #Avg_Cmt_Kj indicate a statistically significant difference at a 95% confidence level (df = 22, p-value = 0.016). Such results support the alternative hypothesis that #Avg_Cmt_Kj is significantly greater than #Avg_Cmt_K. However, the t-test results indicate no statistically significant difference between #Avg_LOC_K and #Avg_LOC_Kj (df = 22, p-value = 0.125). We conjecture this is because Kotlin source files participating in the interactions are often committed to modify only a small amount of code related to the facade, such as method calls, field access, parameter changes.
To answer RQ2, the average number of commits and LOC for 23 open-source Kotlin-Java projects indicate that Kotlin source files participating in Kotlin-Java interactions tend to have more commits but not more LOCs.
@@ 4.3. Common Mistakes in Kotlin-Java Interactions
Table 6 presents the data in the process of locating commits involving problematic Kotlin-Java interactions. Initially, there were 83,792 commits among these projects. After filtering commits with comments containing the words ‘fix’, ‘bug’, ‘issue’ or ‘exception’, there were 5,137 remaining commits. These 5,137 commits were compared with the dependency graphs of the latest version, resulting in 1,314 commits with modified files participating in Kotlin-Java interactions. Finally, we manually inspected these 1,314 commits and confirmed 103 commits containing fixes for problematic Kotlin-Java interactions. After discussion among three developers, we classified 98 of these 103 problematic Kotlin-Java interactions into 7 categories according to the issue and fixing strategies. Table 7 lists these categories and their occurrences. In the following, we provide an example for each of these 7 categories.